Network Status

News & Information

There are no Open Network Issues Currently

Server Status

Below is a real-time overview of our servers where you can check if there's any known issues.

Server Name HTTP FTP POP3 PHP Info Server Load Uptime
Server2 PHP Info
Email Deliverability (Resolved) High

Affecting Server - Server2

  • 08/06/2022 13:40 - 08/07/2022 00:46
  • Last Updated 08/07/2022 00:46

We are aware of an email issue within Server 2 where emails are not getting delivered. We are looking into this issue and hope to have it resolved soon.

 

08/07/2022 0046: Emails have now been restored.

Cron Job (Resolved) Critical

Affecting Server - Server2

  • 03/16/2021 13:32 - 03/20/2021 11:29
  • Last Updated 03/16/2021 17:56

We have experienced an issue with our cron job and our server has terminated some client's accounts that the system believed were overdue. We have interrupted the cronjob process and have begun restoring accounts. We will have further updates as soon as possible.

Update 1: Our billing center has canceled orders, added late fees, and canceled active addons. If you are affected by late fees or your service is canceled, please submit a support ticket and we will take care of your issue as soon as possible. We have already gone through and removed most of the late fees that had been added by the system. 

Update 2: Our server is still in progress with restoring the accounts that have been erroneously terminated.

Update 3: Accounts have been restored. If you have been affected by invoices not being generated, please let us know via support ticket to the billing department and we will take care of your billing issues.

Mail Deliverability (Resolved) Critical

Affecting Server - Server2

  • 08/28/2020 00:00
  • Last Updated 09/08/2020 18:37

We are currently investigating a mail deliverability problem with incoming emails to the server. Mail can still be sent but can not be received. We are actively investigating this issue and will release more information once it becomes available.

Server Migration (Resolved) Medium

Affecting Server - Server2

  • 08/26/2020 21:30 - 09/08/2020 18:36
  • Last Updated 08/27/2020 00:16

Dear Users,

USHost247 will begin migrating from older servers to newer, higher capacity servers. Servers are currently being configured and setup to get ready for the migration of accounts. We will release a new announcement with the updated information for those running custom nameservers. For those that currently use NS1.NOC10.NET and NS2.NOC10.NET as your nameservers, there will be no modifications needed from your end. As with all migrations, there is a chance for error and delays in DNS propagation, which we will keep the old server up and running until the migration is fully complete. We will release more information as it comes available.

RAID Resync (Resolved) Medium

Affecting Server - Server2

  • 02/24/2019 11:34 - 02/26/2019 07:55
  • Last Updated 02/24/2019 11:35

During our monitoring process, we could see that software raid resync going on in the server server2.noc10.net. This would be the server's response to handle a disk issue that happened recently and should be completed successfully without any issues. However, you may notice a slightly degraded server performance during this resync process, until it gets completed.

White Screen During cPanel Login (Resolved) High

Affecting System - cPanel

  • 11/09/2017 21:32
  • Last Updated 11/10/2017 02:05

Users seem to be having a problem logging into cPanel. Once credentials are verified, cPanel loads into a solid white page. 
We are currently looking into this problem and will update this as more information when found out.

>>
Problem has been replicated. Investigating

>>
Problem has been solved.

FTP Connectivity Issues (Resolved) Medium

Affecting Server - Server2

  • 08/24/2017 03:23 - 08/29/2017 10:47
  • Last Updated 08/24/2017 05:45

Users have reported FTP connection issues after server migration. 
We are working on investigating and resolving the issue.

Cron Jobs (Resolved) Medium

Affecting Server - Server2

  • 08/02/2016 08:13 - 08/03/2016 04:47
  • Last Updated 08/02/2016 13:03

We have been informed of a cron job problem that is currently affecting more than one account. We are currently investigating this issue. We will update with any information as it comes available.

*We have identified some problems in the error logs and are currently working to correct these errors and then test crons to make sure they are working correctly. If you use cron jobs in cPanel, please pay attention to make sure they are running correctly. If not, please contact us and let us know.

Datacenter Outage (Resolved) Critical

Affecting Server - Server2

  • 11/02/2015 17:11 - 11/03/2015 12:09
  • Last Updated 11/04/2015 13:57

The links between Newark and our datacenter of Beauharnois are down:
6x100G
20x10G
We will update this taso as soon as we have more information.

There has been a main fiber line cut near the datacenter and has knocked out connectivity to North American BHS datacenter.

Here is the update from OVH:

 

Hello, 

On Monday November 2, 2015, the BHS datacenter suffered a new network problem impacting all BHS Datacenter customers. This incident was similar to 5 months ago, when 3 pairs of our optical fibers were cut between Beauharnois and Montreal on the North route. 

Two years ago, we began adding redundancy on the 103 KMs of fiber on the South loop. Since then, there have been long delays concerning this project, bringing the time of completion to 24 months instead of the planned 12. Work was finished this week and only in response to the urgency of the situation. We forced the installation of the new South fiber optic line, thereby reducing the duration of downtime of the public network from 10 hours and 51 minutes to 4 hours and 55 minutes. 

Now, the datacenter is connected in total redundancy by the North loop and the South loop. We have 5 directions: 2 towards New York, 1 towards Chicago and 2 towards Montreal.

Three years after starting the project, we are continuing the work on improving the 2nd route towards New York where all that remains is to solder the optic fiber at the Canada/USA border. Negotiations are underway to connect Toronto and Detroit, MI.

Here is what our dark fiber and DWDM network connecting BHS looks like: 

We own: 
- 4x10 Gbps between BHS and Montreal via the North loop 
- 4x10 Gbps between BHS and Montreal via the South loop 
- 6x100 Gbps towards New York/Newark via the South loop 
We rent: 
- 10x10 Gbps towards Chicago via South loop 
- 20x10 Gbps towards Newark via the North loop 
- 4x10 Gbps towards London (through Canada) via South loop 
- 4x10 Gbps towards Paris (through USA) via the North loop 

 
 
 
 
 
 


Within 3 to 4 weeks, we will finish the work on our own 2nd route towards New York/Newark and we are going to activate an optical failover of 6x100 Gbps through the following 2 routes: BHS/NYC-NWK. This will allow us to terminate the rental of the 20x10 Gbps through Newark.

 
 
 
 
 
 


We are going to increase the capacity of Montreal/Chicago to 2x100 Gbps then 4x100 Gbps while waiting for our own route to be finished between BHS and Chicago by Toronto/Detroit. We cannot yet provide an ETA, knowing that we lack 800 KMs between Toronto and Detroit, MI. We already have in place BHS towards Toronto and Detroit, MI towards Chicago, IL.

 
 
 
 
 
 

At the same time, we are working on the network in the USA between New York and Newark (NY) then Chicago (IL) and Ashburn (VA). This new USA network is planned to be deployed in Q2 2016 and will not only provide redundancy from BHS towards Chicago through New York but also allow for the connection of a datacenter in the USA. 

 
 

We sincerely apologize for the latest BHS downtime. We underestimated the time and the complexity of the fiber optic network construction projects in Canada, thus creating an unreasonable delay in securing each kilometer connecting the BHS Datacenter. Even though today it's done and we have 100% redundancy, we are really sad to announce this under such circumstance.

Within the next 10 days, we will offer compensation which conforms to our SLA for the month of November.

The exact times concerning the incident: 
- down at 11:53 AM GMT-5 (5:53 PM GMT+1)
- at 4:48 PM GMT-5 (10:48 PM GMT+1), traffic to the BHS datacenter from the outside returned to normal thanks to the optical fiber by the South
- in contrast, the links between BHS and the other OVH datacenters and the VAC (Anti-DDoS) were affected until the complete repair of the North fiber. Traffic was able to resume at 10:44PM (04h44 GMT+1).

Best Regards,
Octave 

Server Restart (Resolved) Medium

Affecting Server - Server2

  • 05/05/2015 22:30 - 05/06/2015 06:22
  • Last Updated 05/05/2015 11:20

We have sceduled a server2 restart at 22:30. This restart will simply install a few updates that have been performed on the server. There will be minimal downtime during the restart. 1-5 minutes.

Directory Issue (Resolved) Critical

Affecting Server - Server2

  • 06/17/2014 18:22 - 06/20/2014 20:35
  • Last Updated 06/17/2014 18:23

We are currently having a directory issue causing user's websites to show a 404 error. We are in progress of getting each account back online. Some users may already have their accounts back online and may have not noticed a difference. Some may take longer than others. No data loss has occurred.

Email (Exim) Server Issues (Resolved) High

Affecting Server - Server2

  • 01/25/2014 05:00 - 01/26/2014 19:31
  • Last Updated 01/25/2014 05:02

We are currently working with the datacenter to get our email server back online. There is a small backlog of queued emails, the emails will get processed as soon as the issue is worked out with the datacenter.

Blank Databases (Resolved) Critical

Affecting Server - Server2

  • 12/05/2013 21:11 - 12/17/2013 19:20
  • Last Updated 12/05/2013 21:12

We are aware of some users having blank databases after the transfer. We have taken backups of the databases from the old server and currently reimporting every database. This will take some time to upload and import. We apologize for any inconveinences. 

Loading cPanel (Resolved) Critical

Affecting Server - Server2

  • 07/07/2013 20:35 - 07/07/2013 21:03
  • Last Updated 07/07/2013 20:38

We have reported an issue to cPanel technicians that has disallowed users access to cPanel and WHM. When trying to load into the login screen, the page will take a long time to load or often time out instead. We still have direct access to the server. We will update everyone with more information as it comes.

cPanel Login/Template Issue (Resolved) High

Affecting System - cPanel

  • 04/11/2013 18:37 - 04/12/2013 00:19
  • Last Updated 04/11/2013 18:38

Clients are having trouble logging into cPanel. Template in cPanel does not display correctly and soon as a user clicks on something, the user is immediately logged out and directed straight to cPanel login screen.

Server Reboot (Resolved) Low

Affecting Server - Server2

  • 02/09/2013 23:59 - 04/11/2013 22:15
  • Last Updated 02/09/2013 03:18

We have a server reboot scheduled for 12am, 2/9/2013. 


CloudLinux has been installed to increase server stability and reliability and a server reboot is required to finish the setup. 

Sites may be inaccessible for 1-2 minutes while the server reboots at 12am.

Network Issue (Resolved) Critical

Affecting Server - Server2

  • 01/15/2013 23:44 - 01/15/2013 03:23
  • Last Updated 01/15/2013 00:03

Websites are going down for various times and coming back online at random times. We had a problem this morning that was said to be resolved; same problem exists. 


Problem has been reported to datacenter and technicians reported back that it is related to a network issue and everyone on the network (including provider itself is affected). Datacenter has been notified of this issue prier to this announcement and they are already working on the issue.

Once we have more information on the issue, we will release in this Network Announcement.

Outage 1/14/13 (Resolved) Critical

Affecting System - Entire Network

  • 01/14/2013 00:00 - 01/14/2013 00:00
  • Last Updated 01/14/2013 17:57

Dear Users,


USHost247 had an outage earlier this morning. After investigation, we had a reply from our datacenter.

We are aware of a network issue this morning that was causing access issues across our entire network.  We have identified the problem and the network is almost back to normal at this time.  We apologize for the inconvenience.

The USHost247 server did not crash or shutdown, this was just a network issue that we had no control over at that point in time. We will release more information if it comes available.

Best Regards,
USHost247 Staff

httpd Service Crash (Resolved) Critical

Affecting Server - Server2

  • 10/06/2012 08:25 - 10/08/2012 19:23
  • Last Updated 10/06/2012 08:29

Lately our HTTPD service has been crashing and restarting itself. Usually ending up 10-20 seconds downtime. (As you may see on HostTracker)


This morning, our httpd service went down for 1 hour. We restarted the service manually and the service came back online. We are looking over all error logs to see what the cause of the crash is. 

No data was lost. No accounts were lost.
(HTTPD is the service that delivers the content to the browser from the server through Apache.)

Best Regards,
USHost247 Staff

Emergency Server Takedown (Resolved) Critical

Affecting Server - Server2

  • 09/17/2012 21:40 - 09/17/2012 22:34
  • Last Updated 09/17/2012 21:31

We will be taking the server down for critical maintenance. Once finished, the server will be back online. Should not take more then 20 minutes.